42. (Problem modified from textbook) A typical adult has an average IQ score of 105 with a standard deviation of 20. If 20 randomly selected adults are given an IQ test, what is the probability that the sample mean scores will be between 95 and 120 points? Round answer to three decimal places (i.e. 0.123) _________Answer
Answer #1

42)

Solution :

= / n = 20 / 20

= P[(95 - 105) / 20 / 20 < ( - ) / < (120 - 105) / 20 / 20)]

= P(-2.24 < Z < 3.35)

= P(Z < 3.35) - P(Z < -2.24)

= 0.9996 - 0.0125

= 0.987